OpenStudy (anonymous):

The endpoints of a diagonal of a square are (5,-1) and (-7,3). Find the coordinates of the two other vertices. A little help please? :)

OpenStudy (anonymous):

np. First get the distance between the 2 points. That distance will be the same for the other two points. For the original 2 points, get slope, then get negative reciprocal for perpendicularity. Get midpoint of original points. For last part, get two points for which this midpoint is also midpoint of the new perpendicular line. This is a highly condensed explanation and we now need to expand this and take it step-by-step.

OpenStudy (anonymous):

Do you know how to get the distance between the 2 original points?

OpenStudy (anonymous):

using the distance formula ^_^

OpenStudy (anonymous):

Good! Can you get the slope?

OpenStudy (anonymous):

Yeah. m = y-y1/x-x1 :)

OpenStudy (anonymous):

If you forgot, it's \[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} } = m\] Okay, looks like you knows that, Great!

OpenStudy (anonymous):

The diagonal of the other points will be perpendicular, so get a new slope for the upcoming next 2 points for that line.

OpenStudy (anonymous):

Negative reciprocal, but I would guess that you know that, too.

OpenStudy (anonymous):

ahh, negative reciprocal because its perpendicular :) yes yes ^^

OpenStudy (anonymous):

With the midpoint of the given 2 points (which will also be the midpoint of the new 2 points) and with the negative reciprocal slope, the new points will be on that line, so you will use the slope formula and midpoint distance twice, once for each new point.

OpenStudy (anonymous):

You might have done so already, but making a drawing will help a bit, if you need to see it and if you have any trouble just visualizing.

OpenStudy (anonymous):

woo. *still trying* ^^

OpenStudy (anonymous):

np. At this point, it's just algebra. I think you have the conceptual underpinings down now.

OpenStudy (anonymous):

haha :) even my classmates are confused. XD

OpenStudy (anonymous):

Another way to have done this problem is to use distance formula, get half of the distance for the radius of a circle with the midpoint of the two points as the center of that circle. Still uses negative reciprocal. It's basically the same approach with a slightly different conceptualization.

OpenStudy (anonymous):

do i have to equate the two equations formed?

OpenStudy (anonymous):

my head my head >< worrrrkkk

OpenStudy (anonymous):

\[y _{1} - y _{midpoint} = (-1/m)(x _{1} - x _{midpoint})\]Yes, equate values, using the above formula rearranged and "half the distance" formula using midpoint and one of the new points.

OpenStudy (anonymous):

Yes, it's a lot of algebra, but it's straight-forward now. I have to believe you can easily follow paper-and-pencil constructions steps to guide your equations. It's a brute-force problem!

OpenStudy (anonymous):

I forgot to ask, you didn't have any trouble getting that midpoint, did you? Do you need help with that?

OpenStudy (anonymous):

(-1,1)?

OpenStudy (anonymous):

Yes, and I can tell that you are fairly advanced and a good student. You're probably taking Analytic Geometry right now?

OpenStudy (anonymous):

Ahm. i think so. haha. i'm only 14

OpenStudy (anonymous):

honestly, it's already 2 am here and i have to make my head work ^^ahahaa

OpenStudy (anonymous):

im sorry if i couldn't get it that fast. :/

OpenStudy (anonymous):

This is an advanced problem for Algebra I. I'm a little surprised that a problem of this complexity was given to you for age 14. Anyway, once you equate the slope formula I wrote out for you as my latest formula and the distance, you basically substitute the right side into the distance formula, so you are getting the value of the x-coordinates of the new points (in the adjusted distance formula). Once you get the x-coordinates, you can go back to the slope formula to get the y's. And no problem on doing this fast or not. It's not really an easy problem. Several steps and a lot of conceptualizing.

OpenStudy (anonymous):

i think i got the one missing vertex. haha! (-3, -5)?

OpenStudy (anonymous):

i also got (1,7) -_-

OpenStudy (anonymous):

@tcarroll010 ? :)

OpenStudy (anonymous):

Still here. As I was posting, I was thinking conceptually and had not done the computations. I would have to get paper and pencil and do them. You want me to check your answers?

OpenStudy (anonymous):

yes yes :)

OpenStudy (anonymous):

For the original 2 points, D = sqrt(160), so d is half which is distance from midpoint to any of the 4 points so d= [sqrt(160)]/2. Midpoint is (-1, 1) which already now by now. The original slope is 4/(-12) so the slope of the new points is m=3. Still typing and figuring.

OpenStudy (anonymous):

I'll be waiting :) haha!

OpenStudy (anonymous):

m = 3 = (y-1)/(x+1), so y-1 = 3/(x+1). d = [(x+1)^2 + (y-1)^2]^(1/2) = [sqrt(160)]/2. Check my algebra for accuracy. Keep me on track.

OpenStudy (anonymous):

Substituting, [sqrt(160)]/2 = [(x+1)^2 + (3/(x+1))^2]^(1/2) Now, checkpoint: did you do this, or at least, does this appear sensible to you at this point?

OpenStudy (anonymous):

sensible :) haha. you gave me ideas and tried and tried

OpenStudy (anonymous):

and I*

OpenStudy (anonymous):

in order to check answers, all sides must have the same distance right? :) :)

OpenStudy (anonymous):

Continuing, squaring both sides: 40 = (x+1)^2 + [3/(x+1)]^2 40(x+1)^2 = (x+1)^4 + 9. Yes, all sides must be same distance, but also, the diagonals distance have to be equal to each other, and that's why we have to go this "long way".

OpenStudy (anonymous):

ohh, i got \[4\sqrt{10}\] as distance of the diagonals ^_^

OpenStudy (anonymous):

Yes, that's what I call D. And d = D/2 = distance from midpoint to any of the four points. Also, I'm pretty confident in the steps I am taking and that last equation I derived which is a quadratic in (x+1)^2. You can substitute n = (x+1)^2 and solve the quadratic in "n", and then once you get "n", you can substitute back for "x". I have to go pick up someone at a hospital and will be gone for a bit. I know it's late where you are, but you have the whole methodology now and most of the equations. You can either wait for me to come back, or you can probably finish this out on your own now. Sorry I have to go.

OpenStudy (anonymous):

Ahm yeah it's okay :) Btw, thank you for your help! :D

OpenStudy (anonymous):

I just got back from the hospital, and [sqrt(160)]/2 = [(x+1)^2 + (3/(x+1))^2]^(1/2) gives x = {1, -3}. Using these in the slope formula will give you your 2 new points (-3, -5) and (1, 7), so you had the right new points. Good job! Sorry for the interruption.

OpenStudy (anonymous):

Ahm again, it's okay :) Thanks again for your help!! :D